OpenStudy (anonymous):

solve the equation 8= c/-4+6 c=

OpenStudy (anonymous):

c = 8*2

OpenStudy (lgbasallote):

\[8 = \frac{c}{-4} +6?\]

OpenStudy (anonymous):

c=8x2?

OpenStudy (lgbasallote):

\[8 = \frac{c}{-4} + 6\] \[8 - 6 = \frac{c}{-4}\] \[2 = \frac{c}{-4}\] can you solve for c now?

OpenStudy (anonymous):

Firstly tell is (-4+6) is in denominator or -4 only in denominator??

OpenStudy (anonymous):

-4 ? only denominator

OpenStudy (anonymous):

Then, c = 2 X (-4) = ???

OpenStudy (anonymous):

-8?

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

Thankyou so much!!!!

OpenStudy (anonymous):

Say thanks to Lgba..

OpenStudy (anonymous):

thanks you too lgbasallote haha